ISS On-Orbit Status 1 Feb 2003


image All ISS systems continue to function nominally, except as noted previously or below.  

On this day of great loss, tragedy and grief, the station crew was given time for quiet mourning after CDR Ken Bowersox was informed of comm loss with Shuttle Columbia at 9:50am EST.  "The cause in which they died will continue.  Our journey into space will go on.‰ (President Bush, 2/1/03).

Work is underway at MCC-Houston to review any potential impacts of the tragedy on ISS program and logistics plans.  No last minute changes to the Progress 10P load manifest were found necessary. Liftoff of the cargo ship at Baikonur is still scheduled for tomorrow morning at 7:59am EST.   Separation command to Progress 9P was given this morning at 10:59am, and successful separation was confirmed at 11:01am.  [ISS attitude, in free drift for the undocking, was then maneuvered to duty attitude and later, at 1:40pm, to experiment attitude to allow observation of the Progress re-entry with the infrared spectrometer of the Russian geophysical GFI-1 "Relaksatsiya" experiment from the Service Module (SM).  Maneuver back to standard attitude occurred at 2:25, followed by handover of control authority to U.S. CMG momentum management.]

Work aboard the station focused on the weekly 3-hr. housecleaning of the station, which includes removal of food waste products, periodic cleaning of compartments with vacuum cleaner, and wet cleaning of surfaces.

The crew also completed the regular daily tasks of SOSh life support systems maintenance (Bowersox) and IMS inventory database auto export/import.

FE-2/SO Don Pettit took two ppCO2 (carbon dioxide partial pressure) readings with the CDMK (CO2 monitoring kit) in SM and Lab, one in the morning, the other before dinner time (2:30pm).

The crew performed their regular 2.5-hr. physical exercise sessions on TVIS (treadmill), CEVIS (cycle ergometer), and RED (resistive exerciser).

Due to the recent trip on the Lab RPCM (remote power controller module), VTR-1 (video tape recorder #1) continues to be unusable in its current configuration.  MCC-H is looking at a potential work around using a specific cable configuration.

Thanks to Don's and Sox's troubleshooting, POC (Payload Operations Center) made good progress with the EXPRESS Rack 2 problem.  Nominal laptop data are being received, and POC is working on a plan to troubleshoot the lack of good Ethernet communications with the RIC (rack interface controller).

A task for familiarizing Don Pettit with MSG (microgravity science glovebox) operations has been added to the "job jar" task list and hard-scheduled for 2/4, prior to the planned MSG ESEM (exchangeable standard electronic module) and PDC (power distribution controller) activities next week.  ESEM and PDC will arrive on Progress 10P on 2/4.

Science Update (Expedition Six -- 9th):

This week  the crew completed the EarthKAM activities. Many students were so excited that it was difficult for some teachers to send them home.  The crew's work on the EarthKAM payload has surely inspired the next generation of scientists and space explorers.  The second InSPACE sample monitor shows that the samples are viable for processing once MSG is functioning. Also, this week the crew completed the second session of Renal Stone collections and a new EVARM activity that will determine a closer correlation between the EVARM badge calibrations and the onboard IV/CPDS in the space environment.

Extra-Vehicular Activity Radiation Monitors (EVARM):  The crew was reminded to return the EVARM badges (4 pouches) to the IV-CPDS (intravehicular charged particle directional spectrometer) location after reading.

GASMAP/Pulmonary Function in Flight (PuFF):  The next PuFF session is planned for Monday, 2/3.

Renal (Kidney) Stone Experiment:  Completed for Increment 6.

Human Research Facility/Workstation (HRF WS):  Continuing.

Ultrasound:  n/a

Foot/Ground Reaction Forces During Space Flight (FOOT):  New run by Bowersox next week.

Pore Formation and Mobility Investigation (PFMI):  Complete for Inc. 5.  On hold pending MSG PDC (power distribution controller) and ESEM3 (exchangeable standard electronic module 3) replacement. Will remain in MSG until further notice.

Space Acceleration Measurement System (SAMS):  SAMS continues to collect acceleration data for vibratory characterization of the microgravity environment.

Microgravity Acceleration Measurement System (MAMS):   MAMS continues measurement of microgravity environment in the quasi-steady regime for general characterization.  

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):   Temperatures are nominal. PCG-STES is currently in heating mode.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   The second performance of the Coil Monitoring Procedure once again provided valuable information that 5 of the 6 Coil Assemblies remain viable for operations and that the particles remain dispersed.  Thanks to the crew for their excellent video and description. Looking forward to initiating operations next week.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.  Was photographed by the crew during the 1/15 EVA.

Zeolite Crystal Growth (ZCG):  ZCG has finished science operations for Inc 6.

EarthKAM (EK):  On behalf of the entire ISS EarthKAM team, as well as the 27 participating middle schools, the ground thanked the crew for supporting their operations this week.  The crew's efforts have helped hundreds of students command over 750 beautiful images. Several teachers have reported that they‚ve had trouble getting their students to go home at the end of the day, since they‚ve been so excited by the images they‚ve been receiving.

Crew Earth Observation (CEO):  The ground has not had such detailed shots of the hard-to-get glaciers on the west side of the Andes as the ISS provided.  These fill a distinct gap in the imagery archive for this cloud-socked part of the world.  Lake Poopo and nearby dry lakes have had an unusual inflow of water (unusual for an inter-ENSO period), presumably related to the very heavy cloud masses we have witnessed for a couple of weeks over this high desert.  The ground is very interested in .WAV files relating to view angles on pack ice and icebergs. Today's targets for the CEO (crew earth observations) program were Bangkok, Thailand (nadir pass; ESC [electronic still camera],) Dar es Salaam, Tanzania(nadir pass; ESC), Patagonian Glaciers (crew was to try for glacier tongues on the west side of the ice fields where glaciers come down to the Pacific Ocean), Congo basin (Dynamic event.  Particularly clear weather as the ITC [Intertropical Convergence] cloud band lies far to the south.  Suggest basic mapping strips along the line of flight for this seldom viewed region.  ISS passed over the vast Congo River [with black forest water], and then the Ubangui River [with a muddy look from the eroding savannas and desert margins that it drains]. Shooting somewhat left into partial sunglint gave the best results for the river patterns that are the major interest), Lagos, Nigeria (nadir view of the vast coastal city; ESC), and Eastern Mediterranean Dust (Sahara dust plumes moving into the Mediterranean across the Egyptian coast were expected as a storm approaches.  Oblique views were best).

ISS Orbit  (as of this morning, 6:39am EST [= epoch]):

  • Mean altitude -- 387.4 km
  • Apogee -- 391.1 km
  • Perigee -- 383.7 km
  • Period -- 92.30 min.
  • Inclination (to Equator) --  51.64 deg
  • Eccentricity -- 0.0005465
  • Orbits per 24-hr. day -- 15.60
  • Mean altitude loss in last 24 hours -- 250 m
  • Solar Beta Angle -- 12.2 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ‚98)  -- 23968
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •  http://www.hq.nasa.gov/osf/station/viewing/issvis.html
  •  

Please follow SpaceRef on Twitter and Like us on Facebook.